# WebVoyager [Uncaptioned image]: Building an End-to-End Web Agent with Large Multimodal Models

### 25 Jan 2024 v1

@@ 3.4 Action Space
We define the action space of our agent similar to how human browse the web. To this end, we implement the most commonly used mouse and keyboard actions, sufficient for the agent to browse various web pages and locate the content required for the task. With the help of numerical labels in screenshots, we enable the agent to respond with a concise Action Format. This approach precisely locates the elements requiring interaction and executes the corresponding actions. The usage and format of actions are presented as follows:
@@
4 Benchmark for WebVoyager
4.1 Website Selection
In this paper, we select 15 popular and representative websites that cover different aspects in our daily life to ensure diversity in our evaluation, including Allrecipes, Amazon, Apple, ArXiv, BBC News, Booking, Cambridge Dictionary, Coursera, ESPN, GitHub, Google Flights, Google Map, Google Search, Huggingface and Wolfram Alpha. Due to technical limitations, we regretfully omit websites that necessitate login or CAPTCHA for accessing their content, but this does not impact the evaluation of effectiveness. Additionally, Google Search is a universal website that can serve as a starting point for any website, making our framework applicable to a wide array of scenarios.
Figure 3: Data creation process using self-instruct.
4.2 Data Construction
We employ a combination of self-instruct (Wang et al., 2022) and human verification to construct our evaluation set. Figure 3 illustrates our data creation process. Initially, we draw inspiration from the web agent tasks in Mind2Web (Yin et al., 2023; Deng et al., 2023), and manually sample and rewrite some tasks for Google Flights, Google Map, Google Search, Booking, and Wolfram Alpha. We prioritize the diversify of tasks such that they can test various functionalities of these websites. This process yields initial seed tasks in the Task Pool for subsequent generation.
In step two, we sample tasks from Task Pool as in-context examples (Dong et al., 2022) and prompt GPT-4 Turbo to generate approximately 100 new tasks (20 iterations). Then we manually verify each generated task and rewrite them if necessary to ensure its high quality and the answers can be found on the corresponding website, then we add them to the Task Pool as additional seed tasks. This step allows us to create human-validated seed tasks for each website.
Finally, the process of step three is very similar to step two, except that we sample more diverse in-context examples in Task Pool and directly add the generated tasks to Task Pool in each iteration. We manually verify that the generated tasks have low repetition and the answers to the generated tasks can be found on the websites. In total, we collected 20 tasks per website, resulting in a total of 300 tasks.
4.3 Annotation Process
After collecting the full task pool, we annotate answers for each task. Since some questions are open-ended and the web information may change, these questions may not have a fixed golden response. Thus we label each data entry with an answer, categorized as either ‘Possible’ or ‘Golden’. For answers labeled as ’Golden’, we provide a comprehensive listing of possible responses and consider them to be stable in a short term. The ’Possible’ category covers the following scenarios: 1) Answers for open-ended tasks where it’s hard to find a exact match answer, such as summarization. 2) There are multiple answers that satisfy the task, making it impractical to list all of them. Therefore, we provide a partial listing. 3) Tasks related to real-time information, where the answer may change frequently, e.g. flight ticket prices. Hence that the ’Possible’ answers are also correct at the time of our experiments. In total, only 22.3% of questions in our benchmark are annotated with golden responses and the rest only has possible answers.
5 Experiment
5.1 Experimental Setup
Dataset and Metrics
We evaluated our agent on the new benchmark introduced in Section 4. Additionally, we extract 90 web browsing tasks (Level 1 and Level 2) from the GAIA dataset (Mialon et al., 2023), which also come with golden responses. The tasks in GAIA are not annotated with specific websites, we instruct the agent to search for answers starting from Google Search. Following WebArena (Zhou et al., 2023), the primary evaluation metric we adopt is the Task Success Rate, measuring the successful completion of tasks, without considering whether the step-by-step execution is optimal. We set the maximum number of interactions for each task to be 15.
Allrecipes Amazon Apple ArXiv GitHub Booking ESPN Coursera
GPT-4 (All Tools)   5 (25%)   3 (15%)   8 (40%)   3 (15%)   9 (45%)   5 (25%)   8 (40%)   8 (40%)
[…]
WebVoyager 12 (60%)   9 (45%) 12 (60%)   9 (45%) 12 (60%)   8 (40%)   9 (45%) 15 (75%)
$*{}^{*}$ 11 (55%) 10 (50%) 12 (60%)   9 (45%) 12 (60%)   7 (35%) 11 (55%) 13 (65%)
Cambridge BBC Google Google Google Huggingface Wolfram Overall
Dictionary News Flights Map Search Alpha
GPT-4 (All Tools)   6 (30%)   2 (10%)   2 (10%) 13 (65%)   11 (55%)   7 (35%)   8 (40%) 32.7%
[…]
WebVoyager 12 (60%) 12 (60%) 10 (50%) 13 (65%)   16 (80%)   9 (45%)   9 (45%) 55.7%
$*{}^{*}$ 16 (80%) 12 (60%) 11 (55%) 12 (60%)   17 (85%) 13 (65%)   9 (45%) 58.3%
[…]
Experimental Details
We employ GPT-4 Turbo with vision (gpt-4-vision-preview) as the backbone model of our agent, which showcases strong semantic and visual understanding capabilities equivalent to GPT-4V (OpenAI, 2023). For baselines, we include the GPT-4 (All Tools), which integrates vision, web browsing, code analysis, and various plugins in one model. Additionally, we consider an text-only baseline where the agent only receives the websites’ accessibility tree as input to prediction actions. For our environment, we used a fixed browser window size of 1024 * 768 pixels, thereby ensuring a consistent size for the screenshots in our observations. We set the temperature to 1 during generation and allows the agent to explore at most 15 steps.
Success Rate Consistency
Agreement $\kappa$
k=1 47.7% 75.3% 0.51
k=2 55.3% 79.7% 0.59
k=3 54.3% 81.3% 0.62
Full 58.3% 85.3% 0.70
5.2 Evaluation Methods
We adopt human evaluation as our main evaluation metric since most of the questions in our benchmark have open-ended answers. In particular, we provide the human evaluators the complete trajectories of the agent’s interaction with the web (all screenshots and all actions), and ask them to provide a binary judgement of whether the agent successfully completed the task. For each task, we ask three expert annotators to provide judgements. The final Task Success Rate reported is the consensus result after discussion, referred to as human labels. We observe that for most tasks, evaluators can determine the correctness of the task based on the final screenshot. For the remaining tasks, the evaluator may require the second-to-last, third-to-last, or even the entire trajectory to make an accurate judgment. On average, it takes an expert about 1 minute to evaluate one task.
Since it’s very expensive to conduct human evaluation for the web agent tasks, we are interseted to see if it’s feasible to leverage an LMM for automatic evalutaion. To this end, we propose to use GPT-4V as an auto evaluator that emulates the behavior of human evaluators to evaluate the navigation trajectories of WebVoyager. In particular, we provide the task, the responses from WebVoyager, and the last $k$ screenshots to the evaluator and ask it to judge whether the agent has successfully completed the task, where $k$ is a hyper-parameter. The prompt of GPT-4V evaluator is shown in Appendix B. We hypothsize that the strong multimodal reasoning ability of GPT-4V allows it to act as a reliable evaluator, which could reduce the burden of human evaluators and significantly lower the cost of future experiments.
5.3 Result
In Figure 4, we present an example to demonstrate how the agent interacts with the Apple website step by step in an online fashion to complete a task. In the final screenshot, the Agent acquires the desired information, then selects the ‘ANSWER’ action to respond and conclude the navigation. This trajectory represents an optimal path without any redundant actions. However, in many cases, there may be some redundant actions in between. Additional examples are provided in the Appendix C to further demonstrate this.
We present the results for our dataset and the extracted GAIA web tasks in Table 1 and Figure 5. We see that WebVoyager outperforms text-only and GPT-4 (All Tools) baselines in the majority of website tasks, while it is slightly lower than Text-only on Cambridge Dictionary and Wolfram Alpha. This is primarily due to the fact that these two websites are more text-heavy than other websites. Since WebVoyager mostly relies on web screenshots for decision making, dense text might not be easy to fully recognize from the image. We think that extracting such text from the HTML to augment the input could be a potential solution to this problem, suggesting a potential direction for future work. In Figure 5, WebVoyager also achieves much stronger performance than both baselines, showing the efficacy of our proposed agent.
Next we delve into the consistency of our evaluations. First we compute the agreement score among our three human annotators before final discussion, we found the Fleiss’s Kappa (Fleiss, 1971) of human annotators to be 0.7, which is substantial agreement. We then report the Agreement (the ratio of overlap) and Kappa ($\kappa$; Cohen 1960) between consolidated human labels and GPT-4V’s judgements in Table 2. Here, $k$ denotes the number of screenshots provided to GPT-4V, with ‘full’ implying the full trajectory. We see that GPT-4V’s agreement with human annotators gradually improves as it receives more information and its final Kappa score also reaches 0.7, which is on par with the agreement among human annotators. The annotation consistency between GPT-4V and humans suggests that using GPT-4V to automatically evaluate multi-modal web agent’s performance is a promising approach and it has great potential for efficient and large-scale applications.
5.4 Discussions
Direct interaction with the websites is necessary From our experience of using GPT-4 (All Tools), the primary limitation of GPT-4 (All Tools)’s performance is rooted in its reliance on Bing search for web browsing, predominantly depending on web pages fetched by Bing. It lacks the capability to directly access certain websites (such as Apple, Amazon, BBC News, etc.) for searching, clicking, or utilizing their sorting functions. This greatly limits the agent’s ability to complete certain types of tasks.
Both text and vision are necessary for generalist web agents. As discussed eariler, WebVoyager struggles with text-heavy websites. On the other hand, we observe that text-only agent demonstrates significantly poorer performance on websites with complex visual elements, such as Booking and Flights, which require interactions with calendars and other intricate components. In these scenarios, the textual input such as accessibility tree becomes highly complex and verbose, making it far less intuitive than using screenshots. Hence it’s necessary to incorporate both modalities of inputs when building the general purpose agents.
Figure 5: Task Success Rate for GAIA Level 1 and Level 2.
Websites with more interact-able elements are more challenging for agents We also calculate the average trajectory length of tasks completed within the maximum number of iterations, as well as the average number of interactive web elements present on the webpage screenshots. Figure 6 illustrates their relationship with the Task Success Rate. We posit that the average trajectory length serves as a measure of a task’s complexity to some extent, while the average number of numerical labels related to the decision-making process reflects the complexity of a webpage. Intuitively, websites depicted in the lower-left corner of Figure 6, characterized by relatively simple webpages and shorter trajectory lengths, are expected to exhibit higher Task Success Rates. As observed in Figure 6, the results largely align with this intuition, though not entirely. It is important to note that our current measure of webpage complexity is somewhat rudimentary.
Figure 6: Factors related to Task success rate: Average number of labels per webpage and Average trajectory length. The color represents the Task success rate of WebVoyager, with darker colors indicating a higher Task success rate.
5.5 Error Analysis
In this section, we discuss and summarize the primary issues encountered by WebVoyager in the task completion process. These challenges will serve as critical entry points for future enhancements of the Task Success Rate and for devising strategies to obtain an Optimal Trajectory. We attribute the failure of the task to the following 4 issues, labeling each failed task with a major issue and showing the ratio in table 3. In Appendix D, we also provide specific examples for each issue.
Navigation Stuck
The most common failure modes for our agent is running out of steps before completing the task. Upon inpsection, we found the following scenarios: 1) If the agent’s search query is not precise and explicit enough, it be overwhelmed by the irrelevant search results. The agent may prefer to browse different results or wait on these incorrect outcomes rather than correct its previous action; 2) When the scroll-able area is only a small part of the screen, the agent might be unable to locate the correct scrolling area, and repeatedly request the execution of useless scrolling actions; 3) Sometimes in the middle of the page, the agent has trouble deciding whether to scroll up or down. Additionally, the agent also has tendency to repeat its previous mistakes due to the input clipping as mentioned in section 3.2. These meaningless or repetitive actions may hinder the completion of the task.
Visual Grounding Issue
The visual grounding ability of our agent still has a large room for improvement. We have observed the following situations: 1) The agent cannot interpret less frequently observed patterns, such as misidentifying characters representing the pronunciations or math formulas; 2) The agent doesn’t recognize the subtle difference between two observations and thinks the execution has failed; 3) The agent selected the wrong element for action execution due to proximity. For example, the confusion between adjacent elements and the misinterpretation of numbers on a calendar as numerical labels. Sometimes textual information plays a significant role, offering valuable cues and assisting in distinguishing between overly dense web elements. We find that incorporating the text content included in Web Elements can alleviate these problems to some extent. However, stronger visual encoder or additional text inputs might be needed to fully address the issue.
Main reasons for
Failure
Ratio
Navigation Stuck 44.4%
Visual Grounding Issue 24.8%
Hallucination 21.8%
Prompt Misalignment 9.0%
Table 3: Main reasons for task failure. Each task is labeled with a primary issue and the ratio of occurrence for each issue is calculated.
Hallucination
Agents sometimes produce seemingly correct answers, which may require checking carefully to identify errors. We observed the following two scenarios: 1) The agent may overlook certain requirements of the task and settle for an answer that is only partially correct. For instance, when asked for the cheapest product, the agent might respond with a cheap product visible in a screenshot, neglecting the need to sort the options first. 2) The agent might execute an seemingly correct action without raising any errors, which actually deviate it from the correct reasoning path. For example, inputting content to the wrong textbox when there are many textboxes in the webpage is still a valid action, yet it would guide the agent to obtain a wrong answer.
Prompt Misalignment
Understanding and following complex prompts, as illustrated in Figure 7, often poses significant challenges. Moreover, longer trajectories may result in excessively lengthy contexts, hindering effective instruction following. Although many of the errors in Navigation Stuck and Hallucination categories can also be attributed to prompt design, we use Prompt Misalignment to categorize the following situations: 1) the agent failed to generate outputs that can be parsed into executable actions, e.g. providing only the ’Thought’ without the corresponding ’Action’; 2) Prematurely terminating the process using the ANSWER action, though agent knows that the task is not yet complete (explicitly mentioned in its answer).
6 Conclusion
In this paper, we introduce WebVoyager, an LMM-powered web agent designed for end-to-end web task resolution. To evaluate the performance of WebVoyager, we collect a variety of web tasks related to 15 websites. We propose Human Evaluation and GPT-4V based Automatic Evaluation, ensuring result reliability with consistent analyses. The results highlight WebVoyager’s potential with a 55.7% Task Success Rate, yet there remains significant room for improvement. A comprehensive Error Analysis of the weaknesses of WebVoyager is provided after the experiments. For future work, we suggest the exploration of better integration methods for visual information and textual information. Additionally, it’s worth investigating in building mutli-modal web agents similar to WebVoyager using open-sourced LMMs.
Limitations
We recognize the follow limitations of our work: firstly, we haven’t support all possible actions in our environment compared to actions a human user might use when browsing the web. e.g. the Drag action on webpages. Supporting such an action is challenging since the degree of Drag is not a finite set. We may allow it to choose the pixel values to be dragged if the Visual Grounding capabilities of LMMs are further enhanced. Secondly, our agent can only analyze PDF files via OpenAI’s Assistant API and doesn’t support other file formats, especially videos. Supporting other file formats would also require non-trivial amount of effort and we consider it to be an important step towards building AGI-like agents.
Appendix A Prompt for WebVoyager
@@
Appendix B Prompt for Auto Evaluation

### 29 Feb 2024 v3

@@ 3.4 Action Space
We define the action space of our agent similar to how human browse the web. To this end, we implement the most commonly used mouse and keyboard actions, sufficient for the agent to browse various web pages and locate the content required for the task. With the help of numerical labels in screenshots, we enable the agent to respond with a concise Action Format. This approach precisely locates the elements requiring interaction and executes the corresponding actions. The usage of actions is as follows (more details in Appendix C): 1) Click. This action involves clicking on an element within a webpage, typically a link or a button. 2) Input. This composite action involves selecting a text box, deleting any existing content within it, and then inputting new content. 3) Scroll. Scrolling is a common operation for browsing webpages, usually involving the vertical movement of the entire page. 4) Wait. Action execution requires time, and this action is often used to wait for web pages to load. 5) Back. This action is used to return to the previous page. 6) Jump to Search Engine. There are often situations where agents get stuck at a certain website without finding an answer. This action enables the agent to jump to a search engine and start anew. 7) Answer. Once all questions in the task are resolved, this action concludes the iteration and provides an answer in line with the task requirements.
4 Benchmark for WebVoyager
4.1 Website Selection
We select 15 representative websites that cover different aspects of our daily life to ensure diversity in our evaluation, including Allrecipes, Amazon, Apple, ArXiv, BBC News, Booking, Cambridge Dictionary, Coursera, ESPN, GitHub, Google Flights, Google Map, Google Search, Huggingface, and Wolfram Alpha. Due to technical limitations, we regretfully omit websites requiring login or CAPTCHA to access their content. Additionally, Google Search is a universal website that can serve as a starting point for any website, making our framework applicable to various scenarios.
Figure 3: Data creation process using self-instruct.
4.2 Data Construction
We employ a combination of self-instruct (Wang et al., 2022) and human verification to construct our evaluation set. Figure 3 illustrates our data creation process. Initially, we manually sample and rewrite some tasks from Mind2Web (Yin et al., 2023; Deng et al., 2023) for websites including Google Flights, Google Map, Google Search, Booking, and Wolfram Alpha. This process yields initial seed tasks in the Task Pool for subsequent generation. In step two, we sample tasks from Task Pool as in-context examples (Dong et al., 2022) and prompt GPT-4 Turbo to generate approximately 100 new tasks (20 iterations). Then we manually verify each generated task and rewrite them if necessary to ensure its high quality and the answers can be found on the corresponding website, then we add them to the Task Pool as additional seed tasks. This step allows us to create human-validated seed tasks for each website. Finally, in step three, we sample more diverse in-context examples in the Task Pool and directly add the generated tasks to the Task Pool in each iteration. We manually verify that the generated tasks have low repetition, and the answers to the generated tasks can be found on the websites. In total, we collected 40+ tasks per website, resulting in a total of 643 tasks.
4.3 Annotation Process
After collecting the full task pool, we annotate answers for each task. Since some questions are open-ended and the web information may change, these questions may not have a fixed golden response. Thus, we label each data entry with an answer, categorized as “Possible” or “Golden.” For answers labeled as “Golden,” we provide a comprehensive listing of possible responses and consider them stable in the short term. The “Possible” category covers the following scenarios: 1) Answers for open-ended tasks where it’s hard to find an exact match answer, such as summarization. 2) multiple answers satisfy the task, making it impractical to list all of them. Therefore, we provide a partial listing. 3) Tasks related to real-time information, where the answer might change, e.g., flight ticket prices. Hence, the “Possible” answers were also correct during our experiments. In total, 22.3% of questions are annotated with golden responses, and the rest only have possible answers.
5 Experiment
Dataset and Metrics
Besides the benchmark introduced in Section 4, we also evaluated on 90 web browsing tasks (Level 1 and Level 2) from the GAIA dataset (Mialon et al., 2023), which also come with golden responses. Since GAIA does not provide specific websites, we instruct the agent to start with Google Search. We further evaluated our agent on the 50 tasks used in SeeAct agent’s online evaluation Zheng et al. (2024) and compared with their results. Following WebArena (Zhou et al., 2023), the primary evaluation metric we adopt is the Task Success Rate, measuring the successful completion of tasks without considering whether the steps are optimal.
Experimental Details
We employ GPT-4 Turbo with vision (gpt-4-vision-preview) as the backbone model of our agent, which showcases strong semantic and visual understanding capabilities equivalent to GPT-4V (OpenAI, 2023). We include the GPT-4 (All Tools) for baselines, which integrates vision, web browsing, code analysis, and various plugins in one model. Additionally, we consider a text-only baseline where the agent only receives the websites’ accessibility tree as input to prediction actions. We used a fixed browser window size of 1024 * 768 pixels for our environment, ensuring a consistent size for the screenshots in our observations. We set the temperature to 1 during generation and allow the agent to explore at most 15 steps.
5.1 Evaluation Methods
We adopt human evaluation as our main evaluation metric since most of the questions in our benchmark have open-ended answers. In particular, we provide the human evaluators the complete trajectories of the agent’s interaction with the web (all screenshots and all actions), and ask them to provide a binary judgement of whether the agent successfully completed the task. For a subset of 300 tasks, we invite three annotators to judge each trajectory to understand the agreement among human annotators.
Even though human evaluations are accurate, they are often not scalable. Hence, we want to see if leveraging a LMM for automatic evaluation is feasible. To this end, we propose to use GPT-4V as an auto-evaluator that emulates the behavior of human evaluators to evaluate the navigation trajectories of WebVoyager. In particular, we provide the task, the responses from WebVoyager, and the last $k$ screenshots to the evaluator and ask it to judge whether the agent has completed the task, where $k$ is a hyper-parameter. The prompt of the GPT-4V evaluator is shown in Appendix B.
Allrecipes Amazon Apple ArXiv GitHub Booking ESPN Coursera
GPT-4 (All Tools)   11.1%   17.1%   44.2%   14.0%   48.8%   22.7%   31.8%   31.0%
[…]
WebVoyager   53.3%   58.5%   65.1%   51.2%   63.4%   43.2%   38.6%   73.8%
[…]
$*{}^{*}$   51.1%±${}_{\pm 2.2\%}$    52.9%±${}_{\pm 1.4\%}$   62.8%±${}_{\pm 2.3\%}$   52.0%±${}_{\pm 1.3\%}$   59.3%±${}_{\pm 3.7\%}$   32.6%±${}_{\pm 2.7\%}$    47.0%±${}_{\pm 1.3\%}$    57.9%±${}_{\pm 2.7\%}$
Cambridge BBC Google Google Google Huggingface Wolfram Overall
Dictionary News Flights Map Search Alpha
GPT-4 (All Tools)   25.6%   9.5%   2.4%   53.7%   60.5%   37.2%   52.2%   30.8%
[…]
WebVoyager    65.1%   61.9%   59.5%   70.7%   76.7%   44.2%   63.0%   59.1%
[…]
$*{}^{*}$   71.3%±${}_{\pm 1.3\%}$    60.3%±${}_{\pm 2.8\%}$   51.6%±${}_{\pm 1.4\%}$   64.3%±${}_{\pm 2.8\%}$   77.5%±${}_{\pm 2.7\%}$   55.8%±${}_{\pm 2.3\%}$    60.9%±${}_{\pm 2.2\%}$    57.1%±${}_{\pm 0.2\%}$
[…]
Success Rate Consistency
Agreement $\kappa$
k=1 47.7% 75.3% 0.51
k=2 55.3% 79.7% 0.59
k=3 54.3% 81.3% 0.62
Full 58.3% 85.3% 0.70
5.2 Result
Figure 4 presents an example that demonstrates how the agent interacts with the Apple website step by step in an online fashion to complete a task. In the final screenshot, the Agent acquires the desired information, then selects the “ANSWER” action to respond and conclude the navigation. Additional examples are provided in the Appendix D.
We present the results for our dataset and the extracted GAIA web tasks in Table 1 and Figure 5. WebVoyager outperforms text-only and GPT-4 (All Tools) baselines by large margins in most website tasks, while it is slightly lower than Text-only on Allrecipes and similar to Text-only on Github, ESPN, Cambridge Dictionary and Wolfram Alpha. This is primarily because these websites are more text-heavy than others. Since WebVoyager mostly relies on web screenshots for decision-making, dense text might not be easily recognizable from the image. We think extracting such text from the HTML to augment the input could be a potential solution to this problem, suggesting a direction for future work. In Figure 5, WebVoyager also achieves much stronger performance than both baselines. Finally, WebVoyager has a success rate of 30% on the SeeAct online test set whereas the best SeeAct autonomous agent has 26%, showing the efficacy of our proposed agent.
We report the Agreement (the ratio of overlap) and Kappa ($\kappa$; Cohen 1960) between consolidated human labels666the Fleiss’s Kappa (Fleiss, 1971) of human annotators before any discussion is 0.7, which is substantial agreement. and GPT-4V’s judgments on the subset of 300 tasks in Table 2. Here, $k$ denotes the number of screenshots provided to GPT-4V, with “full” implying the full trajectory. GPT-4V’s agreement with human annotators gradually improves as it receives more information, and its final Kappa score also reaches 0.7, which is on par with the agreement among human annotators. The consistency between GPT-4V and humans suggests that GPT-4V is a promising automatic evaluator for multi-modal web agents.
5.3 Discussions
Direct interaction with the websites is necessary From our experience of using GPT-4 (All Tools), the primary limitation of GPT-4 (All Tools)’s performance is rooted in its reliance on Bing search for web browsing, predominantly depending on web pages fetched by Bing. It cannot directly access certain websites (such as Apple, Amazon, BBC News, etc.) for searching, clicking, or utilizing their sorting functions. This greatly limits the agent’s ability to complete certain types of tasks.
Both text and vision are necessary for generalist web agents. As discussed earlier, WebVoyager struggles with text-heavy websites. On the other hand, we observe that text-only agent demonstrates significantly poorer performance on websites with complex visual elements, such as Booking and Flights, which require interactions with calendars and other intricate components. In these scenarios, the textual input such as accessibility tree becomes highly complex and verbose, making it far less intuitive than using screenshots. Hence it’s necessary to incorporate both modalities of inputs when building the general purpose agents.
Figure 5: Success Rate for GAIA Level 1 and Level 2.
Websites with more interactable elements are more challenging for agents. We also calculate the average trajectory length of tasks completed within the maximum number of iterations, as well as the average number of interactive web elements present on the webpage screenshots. Figure 6 illustrates their relationship with the Task Success Rate. We posit that the average trajectory length serves as a measure of a task’s complexity to some extent, while the average number of numerical labels related to the decision-making process reflects the complexity of a webpage. Intuitively, websites depicted in the lower-left corner of Figure 6, characterized by relatively simple webpages and shorter trajectory lengths, are expected to exhibit higher Task Success Rates. As observed in Figure 6, the results largely align with this intuition.
Figure 6: Factors related to task success rate. We show the average number of elements per page and the average trajectory length for each website type. The darker colors indicate a higher task success rate.
5.4 Error Analysis
In this section, we discuss and summarize the primary issues encountered by WebVoyager in the task completion process. These challenges will serve as critical entry points for future enhancements of the Task Success Rate and for devising strategies to obtain an Optimal Trajectory. We sampled 300 tasks from our benchmark and manually labeled the error category for each failed case, we show the distribution of errors in table 3. In Appendix F, we also provide specific examples for each issue.
Navigation Stuck
The most common failure is running out of steps before completing the task. There are three failure scenarios: 1) When the agent’s search query is not precise and explicit enough, it will be overwhelmed by irrelevant search results. The agent may prefer to browse different results or wait for incorrect outcomes rather than correct its previous action; 2) When the scroll-able area is very small, the agent might not be able to locate the correct scrolling area and repeatedly request the execution of useless scrolling actions; 3) Sometimes in the middle of the page, the agent has trouble deciding whether to scroll up or down. The agent also tends to repeat its previous mistakes due to the input clipping, as mentioned in section 3.2. These meaningless or repetitive actions may hinder the completion of the task.
Visual Grounding Issue
The visual grounding ability of our agent still has a large room for improvement. We observe the following issues: 1) The agent cannot interpret less frequently observed patterns, such as misidentifying characters representing the pronunciations or math formulas; 2) The agent doesn’t recognize the subtle difference between two observations and thinks the execution has failed; 3) The agent selects the wrong element for action execution due to proximity. For example, the model sometimes confuses adjacent elements and misinterprets numbers on a calendar as numerical labels. Sometimes textual information plays a significant role, offering valuable cues and assisting in distinguishing between overly dense web elements. We find that incorporating the text content included in Web Elements can alleviate these problems to some extent. However, a stronger visual encoder or additional text inputs might be needed.
Main reasons for
Failure
Ratio
Navigation Stuck 44.4%
Visual Grounding Issue 24.8%
Hallucination 21.8%
Prompt Misalignment 9.0%
Table 3: Distribution of main failure reasons.
Hallucination
Agents sometimes produce seemingly correct answers, which may require checking carefully to identify errors. We mainly see the following two scenarios: 1) The agent may overlook certain task requirements and settle for an answer that is only partially correct. For instance, when asked for the cheapest product, the agent might respond with a cheap product visible in a screenshot, neglecting the need to sort the options first. 2) The agent might execute a seemingly correct action without raising any errors, which deviate it from the correct reasoning path. For example, inputting content to the wrong text box when there are many text boxes on the webpage is still valid, yet it would guide the agent to obtain a wrong answer.
Prompt Misalignment
Understanding and following complex prompts, as illustrated in Figure 7, often poses significant challenges. Moreover, longer trajectories may result in excessively lengthy contexts, hindering effective instruction following. Although many of the errors in Navigation Stuck and Hallucination categories can also be attributed to prompt design, we use Prompt Misalignment to categorize the following situations: 1) the agent fails to generate outputs that can be parsed into executable actions, e.g. providing only the ‘Thought’ without the corresponding ‘Action’; 2) Prematurely terminating the process using the ANSWER action, though agent knows that the task is not yet complete (explicitly mentioned in its answer).
6 Conclusion
We introduce WebVoyager, an innovative web agent powered by large multimodal models (LMM) that can complete real-world web tasks end-to-end by interacting with websites. We have shown through evaluations that WebVoyager outperforms several baselines by leveraging both visual and textual signals. We also propose an automatic evaluation protocol by leveraging GPT-4V as the evaluator for online agents. Our work demonstrates the promise of using advanced LMM capabilities for building intelligent web agents. We hope WebVoyager provides a strong foundation for future research toward building more versatile and capable web assistants.
Limitations
We recognize the following limitations of our work. First, we haven’t supported all possible actions in our environment compared to actions a human user might take when browsing the web. e.g. the Drag action on web pages. Supporting such an action is challenging since the degree of Drag is not a finite set. We may allow it to choose the pixel values to be dragged if the Visual Grounding capabilities of LMMs are further enhanced. Second, our agent currently can only analyze basic file formats (such as text files and PDF files) and doesn’t support all file formats, especially videos. Enhancing support for additional file formats is a crucial step in the development of web agents and we leave it for future work.
Regarding the potential risks of our work, we believe that it requires a substantial amount of safety checks before deploying web agents like WebVoyagar into real-world applications, as the agent might unintentionally download malicious content from unauthorized websites, or input private/confidential information on public websites. Therefore it’s necessary to take extra caution when using and testing our agent.
Appendix A Prompt for WebVoyager
Appendix B Prompt for Auto Evaluation
Appendix C Action Space
We detail the interaction actions that WebVoyager employs to navigate and operate within web environments. These actions are fundamental to how the agent interacts with web pages, retrieves information, and performs specific tasks as part of its operational protocol. The actions range from basic web navigation to more complex operations, tailored to efficiently gather data and respond to queries. Each action is designed with a specific format for easy identification and execution.
@@ Appendix C Action Space
Appendix D Additional Trajectories
Appendix E Additional Related Work
Vision-based Agents Concurrent to our work, a few related works also studied vision-based autonomous agents. VisualWebArena Koh et al. (2024) extends WebArena with additional websites and tasks that focus on visual reasoning to facilitate research on vision-based web agents. SeeClick Cheng et al. (2024) focused on finetuning a LMM to solely leverage screenshots as inputs to interact with websites. WebVLN Chen et al. (2023) introduced a web simulator that provides both HTML text and screenshots to finetune supervised vision-language models. GPT-4V Wonderland Yan et al. (2023) and AppAgent Zhang et al. (2023) instead focuses on building agents that can operate smart phone apps using the GPT-4V as the backbone. These works further underscore the promising prospects in this field.
Large Multimodal Models. In recent years, significant strides have been made in unifying image and text representations within a single multimodal model through joint training with image and text (Li et al., 2019; Dosovitskiy et al., 2020; Wang et al., 2021; Dai et al., 2022; Aghajanyan et al., 2022). Large Multimodal Models (LMMs), following in the footsteps of Large Language Models (Brown et al., 2020; Chen et al., 2021; Chowdhery et al., 2023), attain the capability of instruction following (Ouyang et al., 2022) and exhibit robust multimodal comprehension. Represented by GPT-4V (OpenAI, 2023) and Gemini (Team et al., 2023), LMMs have demonstrated impressive performance on benchmarks (Goyal et al., 2017; Lu et al., 2022; Zellers et al., 2019; Hessel et al., 2022), establishing a foundation for the construction of multimodal agents in subsequent research.